# CataractBot: An LLM-Powered Expert-in-the-Loop Chatbot for Cataract Patients

### v4

@@ 6.1.1. Reasons for (Not) Using CataractBot
Information seekers used CataractBot to address questions they had forgotten or were uncomfortable to ask the doctor or patient coordinator, clarifying answers, and seeking updates. Doctors stated that most information seekers in recent years come prepared for face-to-face conversations, optimizing their limited time with experts and avoiding multiple visits. CataractBot can alleviate the burden of remembering every question. Ten participants highlighted that the bot helped in recalling information provided during in-person consultations, which was either forgotten or incompletely understood. As A13 stated: “We are humans… cannot pay complete attention all the time.” Patients found comfort in “written” information, enabling them to refer to it when needed (similar to (Jain et al., 2018)), “I need not remember everything.” (A20). Six participants also mentioned using the bot to “double-check” (A7) and seek “reassurance” (A17), in spite of knowing certain answers.
@@
Due to existing power differences between doctors and patients (Joseph-Williams et al., 2014), individuals often hesitate to ask questions. Experts acknowledged that educated individuals have access to other information sources, and are comfortable posing questions. In contrast, “less educated patients who don’t speak English” are the ones for whom CataractBot “can do wonders” (D3). Five participants, including three with Grade-12 education, exclusively relied on CataractBot (asking 57 questions in total) as no one in their social circles had experience with cataract surgery, and they were not accustomed to seeking information online. Another expressed that checking with the bot was more comfortable than approaching a doctor, citing the bot’s non-judgemental nature: “(The patient) has cholesterol issues and heart surgery… the doctor would make faces when I ask about food intake… but not the bot.” (A4).
The asynchronous feature was helpful, as the information seekers did not have to constantly wait, either physically or over a call, for an expert’s response. Seven participants emphasized that the bot saved them time, reducing the need to visit or call the hospital. Traveling can be time-consuming, and contacting busy hospitals over the phone is challenging due to high patient loads, as “the phone is always busy” (A4). Information seekers found the bot easy to use, with 86.2% rating ‘Agree’ or ‘Strongly Agree’ in response to the usability question (Appendix A.2), mainly due to their familiarity with WhatsApp.
Experts used CataractBot driven by their commitment to help patients.
“I want to help my patients… I am not bothered about the timing. If they are having a problem, they should be answered… after all, we are living for the patients’ sake.” – D2.
Experts also appreciated the intermediary role of the bot, introducing a layer of separation from patients: “From a doctor’s standpoint, it is very good because it puts in a curtain between us.” (D4).
@@
Despite the compelling reasons to use the bot, out of the 49 information seekers, 8 did not use the bot, sending two or less messages. Six participants mentioned that their in-person interactions with the doctor and patient coordinator sufficed to address their queries.Two participants considered cataract surgery to be a “simple” procedure, leading them to have no questions. This perception stemmed from their previous experience with cataract surgery of their other eye. This suggests that CataractBot could play a more pivotal role in primary healthcare centers, where the human infrastructure may be less capable of addressing information needs or where power imbalances make individuals less likely to seek answers. Additionally, a similar bot could be valuable for more complex (surgical) procedures for which information is less available.
@@
Participants complained about receiving “I don’t know” (IDK) responses (57 total) and experiencing delays while experts verified responses. A total of 36 (10.5%) messages requested status updates—seeking information about the surgery time and discharge time. Due to privacy concerns, CataractBot was not integrated with the hospital’s patient management system, resulting in IDK responses. Patient coordinators later provided answers. While participants acknowledged the appropriateness of IDK responses due to the lack of access to patient details, they suggested providing a “tentative response” if possible. E.g., for the question “Mine is general anaesthesia or local anaesthesia?” (A10), the bot provided an IDK response. However, A10 expected a response such as “99% of cataract surgeries at Sankara Eye Hospital are done under local anesthesia”. From an expert’s perspective, responding to certain IDK questions was challenging as they lacked the full context. E.g., A21 asked: ‘I want to postpone the surgery to February… is this okay or do I need to get this operated immediately?” In response, the doctor corrected the bot’s IDK response with: “Please come and check”. This correction frustrated the attendant, who stated: “But (the patient) already went through the complete eye check; the doctor know her condition… I don’t want to come and visit the doctor [again]… it will waste 2-3 hours.” (A21). For time-sensitive questions, A1 and A9 desired “quick response” from the bot. The delay in expert verification led them to perceive the bot as ineffective.
“The dosage [of eye drops] is 1 drop every 2 hours. If it [bot] takes more than 2 hours to respond [to my eye drop related question], the purpose is defeated” – A1.
6.1.2. Workflow
Out of the total 343 messages sent by patients and attendants, almost all were sent between 7 am and 11 pm, with an average of 14.313.7 messages/hour. The messages were distributed throughout the day, as A18 mentioned: “Whenever I am idle at work or at home… whatever (question) comes to my mind, I just ask.” The workflow employed by participants varied, with time-sensitivity being a crucial factor. A majority of participants (86.4%) agreed that the bot responded quickly. On average, the bot took 11.810.2 seconds to respond. Expert verification took 162.9172.3 minutes. We found verification without edits (151.1147.1 minutes) to be significantly faster than verification with edits (191.3172.4 minutes) with t(265)=2.7, p¡0.05. For time-insensitive questions, participants were “okay to wait” for the expert’s verification, since the response was being verified by “busy doctors”. This demonstrated their clear understanding of the bot’s workflow.
A few participants mentioned that they hardly noticed the wait as they were occupied with their day-to-day activities. They would ask a question, leave their phones, and check the answer later when free. In such cases, notifications played a key role. The message stating that the expert had provided a verified response (Figure 1G), made the otherwise unnoticeable green tick (Figure 1F) prominent and the bot’s process self-explanatory.
@@ 6.1.2. Workflow
If participants received an IDK response, they sought information from other sources, including Google and asking friends and family. For time-sensitive questions, participants reported using Google to “double-check” the initial LLM-generated, unverified response due to the delay in obtaining an expert verified response. Four participants, accustomed to relying on Google for their informational needs, reserved CataractBot exclusively for questions requiring expert opinions, such as “When can i do exercises as anulom vilom?” (P27). According to D4, without CataractBot, patients might have resorted to “Google” and risked following inaccurate advice, given the unavailability of doctors for immediate consultations. CataractBot improves the likelihood of receiving accurate responses, as it uses a custom knowledge base.
We found that the experts were “fine” with the additional task of verifying the bot’s responses, as articulated by D4: “It takes only about 15-20 seconds to answer a question… It was not really eating into my time… It’s much easier than replying to my mailbox.” The asynchronous nature of the bot provided experts the flexibility to offer verifications at their convenience. The majority of responses (31, 10.8%) were provided by experts at 4-5 pm, post their busy workday, as most OPD and surgeries finish by 4 pm. The second-highest responses (28, 9.8%) occurred at 1-2 pm, during their lunch break. This usage pattern was described as:
“I use my mobile phone only when I go for breaks… lunch break, coffee break, or in the morning before I start OPD. The rest of the time, my (cellular) data is off.” – D1.
As five experts used CataractBot on their personal WhatsApp accounts, it minimized the learning curve but blurred the boundary between their home and work, similar to previous findings (Mols and Pridmore, 2021; Wang et al., 2020). Our experts also noted that WhatsApp had a low signal-to-noise ratio, resulting in instances where they missed CataractBot messages: “My WhatsApp has 5000+ [unread] messages… college groups, family groups… I just see messages in the important groups.” (D2). Addressing this issue, the escalation doctor D4 suggested that experts should be recommended to “pin” CataractBot on WhatsApp, ensuring it remains consistently at the top of their chat list.
6.1.3. Question Type and Style
To analyze the medical and logistical questions posed by participants, one of the researchers manually classified them into 12 broad categories, identified in a bottom-up manner (similar to (Li et al., 2023; Clarke et al., 2016b)). Note: Several questions fell into multiple categories. Figure 4A illustrates the relative significance of these categories before, on, and after the day of surgery. Prior to the surgery, patients and attendants were concerned about the surgical procedure (asking questions about “anaesthesia”, “lens”, and “duration of surgery”) and sought guidance on how to prepare for it (e.g., “eat or not before the surgery”, “documents to bring for surgery”, and “take blood thinners on the day of surgery”). One day before the surgery, questions related to the “schedule” peaked, as information seekers awaited calls from patient coordinators about their surgery time. On the day of surgery, questions focused on discharge procedures and post-surgery medication and dosage. After the surgery, participants were concerned about pain management and sought guidance on dos-and-don’ts during recovery (e.g., “washing hair”, “doing yoga”, and “watching TV”).
39 messages from 5 patients and 13 attendants showed signs of anthropomorphizing the bot, addressing it as ‘Doctor’, ‘Sir’, or ‘Ma’am’. Many queries were prefaced with greetings (like ‘hi’) and included acknowledgments (like ‘okay’), contributing to the humanization of the bot. In Indic languages, patients employed honorifics when referring to the bot in their messages and during interviews. Notably, two low-literate patients treated the chat as if they were conversing back-and-forth with a human, providing non-actionable status updates, such as P29 telling CataractBot that “The date of the surgery is tomorrow. We will leave home at 8 am from Ayyappa Nagar.”
6.1.4. Multiple Stakeholders
Patients sent a total of 89 messages (4.75.4 message/patient), while attendants sent 254 messages (8.56.8 message/attendant), significantly more than patients, with t(47)=2.04, p¡0.05 (Figure 4C). Among the 309 verified medical and logistical answers, 62.5% were verified by the expert doctor/coordinator and the rest 37.5% by the escalation expert. This highlights the significance of the bot supporting multiple information seeker and information verifier roles. During our interviews, we identified a few reasons for the patients to minimally use the bot, particularly related to demographics. Patients were typically older with lower levels of literacy and tech-literacy compared to their attendants These patients placed trust in the hospital to provide them with all the required information and did not actively seek information themselves. For instance:
“Dad didn’t try… I was the one thinking about it [the surgery]. He was like, “It’s OK! Not everyone needs to bother in the family… (to know) about what will be this, what will be that”. – A3.
Additionally, patients were advised to minimize screen use post-surgery, contributing to their reduced bot interaction. Interestingly, patients sometimes offloaded the knowledge-seeking ‘work’ to their active attendants—for instance, “I have been using it. I told him [the patient] what (answers) I got. So maybe he did not feel the need (to use it).” (A20)—so the involvement of additional stakeholders had a partially negative impact.
Moving on to experts, both the doctors and patient coordinators highlighted the significant role of escalation experts. Among the medical questions, 76.2% were verified by operating doctors and 23.7% by the escalation doctor. Doctors mentioned having 2-3 days per week reserved for surgeries, and on these “operating days”, they rarely checked their phones during work hours, leading to more escalations. Despite the crucial role of escalation experts, there was a notable difference in their interactions compared to operating doctors and coordinators. While the latter had previous face-to-face meetings with patients and attendants, there was no such guarantee for escalation experts. As a result, escalation experts lacked knowledge of patients’ context, and hence were restricted to providing generic responses.
6.1.5. Question Time and Response Time
A majority of messages (71.4%) were sent in the 8-day period from 5 days prior to surgery to 2 days post-surgery with the highest number of messages (76) sent on the day of the surgery (Figure 4B). 156 messages were sent before the surgery day and 111 post the surgery day. We conducted a one-way ANOVA on the number of messages asked by each patient-attendant group pre-surgery, on the day of surgery, and post-surgery, and found statistical significance with F(2, 60)=3.69, p¡.05. Post-hoc pairwise comparisons with Bonferroni correction indicated a statistically significant difference between ‘pre’ and ‘on’ surgery (t(30)=2.7, p¡.05) and between ‘pre’ and ‘post’ surgery (t(30)=3.2, p¡.05). This suggests that patients and attendants are more inquisitive before surgery, potentially due to heightened anxiety, aligning with prior work(Beisecker and Beisecker, 1990). The higher pre-surgery usage may also be due to the novelty effect. With respect to the types of queries, logistical questions were mainly asked before the surgery, reaching their peak one day before the surgery day (Figure 4B), as anticipated. In contrast, experts expressed that questions on the day of surgery might be unnecessary:
“Multiple questions were asked on the day of surgery, “When is he going to be discharged?”… You are in the hospital… Why use the bot?!… You can walk 10 steps and ask the staff.” – D1.
Regarding the usage patterns of experts, they refrained from using their phones in front of patients due to the negative perception associated with using phones for personal reasons. Opinions on using the chatbot after work hours varied: two experts were reluctant, while four were in favor of using it “anytime”.
6.2. Features
In this section, we address RQ2 by discussing the usage of different features by information seekers and experts.
6.2.1. Experts-in-the-loop
The novel feature of having human experts verify and update the auto-generated responses was appreciated by participants, serving as the key reason for their engagement with the bot. For instance, “That’s why I gave my WhatsApp (number to onboard the bot), so I can talk directly to the doctor.” – A2. Contrary to initial expectations that the icons might be complex, participants found them self-explanatory:
However, since participants knew that these messages were being reviewed by doctors and coordinators, four used the bot to send direct messages to them instead of asking questions (e.g., “Dr. Give me a call?” (P19))
The length of expert corrections in characters varied significantly (t(91)=5.2, p¡0.05) with 152.4104.9 for medical questions and 50.942.6 for logistical questions. Further analysis revealed that their suggested edits fell into five categories: adding missing information (76.8%), asking clarification questions (14.6%), making factual corrections (11.0%), asking the patient to visit the hospital (8.8%), and removing unnecessary information (3.7%).
Finally, experts did not have access to the final answer sent to information seekers after their suggested edits. This limited verification to a single round, minimizing their workload. However, without this transparency, experts were sometimes unsure whether CataractBot had understood their corrections. In the CataractBot system, only the knowledge base update expert had access to the updated answer sent to the information seeker.
6.2.2. Input Modality
f the 343 messages sent by information seekers, 206 (60.1%) were text, 43 (12.5%) audio and 94 (27.4%) related questions. 19 participating patients sent 89 messages (43 text, 26 audio, and 20 taps), while the 30 attendants sent 254 messages (163 text, 17 audio, and 74 taps) (Figure 4C). One-way ANOVAs on message types found statistical significance for attendants (F(2, 87)=13.74, p¡.001), but not for patients (F(2, 54)=0.77, p=0.46). Post-hoc pairwise comparisons, corrected with Holm’s sequential Bonferroni procedure, indicated statistically significant differences between ‘text’ and ‘audio’ (t(29)=4.8, p¡.01) and ‘text’ and ‘taps’ (t(29)=3.0, p¡0.01) for attendants. The minimal usage of audio by attendants could be due to demographic differences, with the majority being under 45 years and well-educated. Audio messages were predominantly used by older semi-literate patients. For instance, “I can’t read (or) write Hindi that well, so asking and listening to audio, I did that.” (P12).
Post-surgery limitations on screen use also made speech input and audio output the preferred modality for patients, with 18 out of 26 audio questions by patients asked post-surgery. 31 participants used the related questions at least once. As P7 stated: “I used the audio message once. After that there was no need for me [to type or use audio]… I just kept selecting from the [related questions] options.” While they found the related questions feature convenient and frequently used it, there were instances of dissatisfaction when (in eight cases) the answers to those resulted in IDK responses. This issue arose because the CataractBot system uses an LLM call to suggest three questions based on the users’ last question, independent of the knowledge base. This approach aimed to uncover gaps in the knowledge base by not restricting questions to the curated knowledge base, and introduce personalization by suggesting questions relevant to the ongoing conversation.
Note: On the expert end, they could only interact with CataractBot using text. This design choice facilitated a “hygiene check” (D4) for thoroughness and precision in experts’ edits, aligning with previous findings (Li et al., 2023).
6.2.3. Multilinguality
Experts interacted with CataractBot only in English. Among our patients and attendants, the majority (29 out of 49) selected English, while 7 opted for Hindi, 6 Kannada, 5 Tamil, and 2 Telugu. None of the participants changed their initially selected language. Given the challenges of typing in Indic languages on smartphones (Joshi et al., 2011), our non-English participants predominantly relied on speech input and selecting from related questions. In total, patients who chose English typed 35 questions, asked 0 questions in speech, and tapped 6 related questions. In contrast, patients opting for Indic languages typed 8 questions, sent 26 audio questions and tapped 14 related questions in total. A similar trend was observed among attendants. Patients found value in local language interaction and speech as input, helping them use the bot independently: “Without even asking me for help, he (patient, his father) was able to use that bot as it supports Tamil… I felt it’s very useful. Because now, for elders… it is very hard to use apps and things.” – A13.
The audio queries in Indic languages were transcribed and translated to English using Azure AI services. However, this was not highly accurate due to the variability in Indic language dialect. This is a known issue and prior work (Jain et al., 2018) relied on human wizards to correct such errors. Experts faced challenges in “deciphering” such messages since they only received the English transcription.
It is common to type Indic languages in Latin script (e.g., instead of using Devanagari script for Hindi), occurring in three messages. This was not supported by CataractBot due to current translation technology limitations, and it responded with “You seem to be asking the question in a language different from the selected language…”. However, when a message contained a code-mix of Hindi and English, the bot responded accurately:
@@ 6.2.3. Multilinguality
Even in the selected language, those with limited literacy struggled to formulate questions (similar to findings in prior research (Jain et al., 2018)). For example, “Tomorrow operation When should come No phone call received yet?” (P18). There were 65 such questions, of which 26 were asked using audio in an Indic language. Despite this, the GPT-4 model’s robustness enabled the chatbot to answer most (45) of such ill-formulated questions appropriately.
6.2.4. Knowledge Base Updation

### v5

@@ 6.1.1. Reasons for (Not) Using CataractBot
Information seekers used CataractBot to address questions they had forgotten or were uncomfortable to ask the doctor or patient coordinator, clarifying answers, and seeking updates. D1 stated that most information seekers come prepared for face-to-face conversations, optimizing their limited time and avoiding multiple visits. CataractBot can alleviate the burden of having to remember every question. Ten information seekers highlighted that the bot helped recall information provided during in-person consultations, which they had either forgotten or not completely understood. Patients found comfort in “written” information, available as a reference (similar to (Jain et al., 2018)), “I need not remember everything.” (A20). Six participants also mentioned using the bot to “double-check” (A7) and seek “reassurance” (A17), despite already knowing certain answers.
@@
Due to existing power differences between doctors and patients (Joseph-Williams et al., 2014) and the judgmental attitudes of some doctors, individuals often hesitate to ask questions. Experts acknowledged that educated individuals are comfortable posing questions. In contrast, “less educated patients who don’t speak English” are the ones for whom CataractBot “can do wonders” (D3). Five participants, all with a Grade-12 education or below, relied heavily on CataractBot asking 57 questions in total, as they reported being unaccustomed to seeking information online. One of them expressed that checking with the bot was more comfortable than approaching a doctor, citing the bot’s non-judgmental nature: “(The patient) has cholesterol issues and heart surgery… the doctor would make faces when I ask about food intake… but not the bot.” (A4).
@@
Despite these reasons, 8 out of 49 information seekers used the bot minimally, sending only one or two messages. Two key reasons emerged. First, six participants mentioned that their in-person interactions with the doctor and patient coordinator sufficed to address their queries. This suggests that CataractBot could play a more pivotal role in primary healthcare centers, where the human infrastructure may be less capable of addressing information needs or where power imbalances make individuals less likely to seek answers. Second, two participants considered cataract surgery to be a “simple” procedure, leading them to have no questions. This indicates that a similar bot could be valuable for more complex (surgical) procedures.
6.1.2. “I don’t know” Responses.
@@
Participants complained about receiving “I don’t know” responses (57 total) A total of 36 (10.5%) messages requested status updates—seeking information about the surgery time and discharge time. Due to privacy concerns, CataractBot was not integrated with the hospital’s patient management system, resulting in “I don’t know” responses. Patient coordinators later provided answers. While participants acknowledged the appropriateness of “I don’t know” responses, they suggested providing a “tentative response” (A10) if possible. For instance, to the question “Mine is general anesthesia or local anesthesia?” A10 expected an answer such as “99% of cataract surgeries at Sankara Eye Hospital are done under local anesthesia”. From an expert’s perspective, responding to certain “I don’t know” questions was challenging as they lacked the full context. E.g., A21 asked: ‘I want to postpone the surgery to Feb… is this okay or do I need to get this operated immediately?”, receiving an “I don’t know” response. To correct this, the doctor said: “Please come and check”. This correction frustrated the attendant, who stated: “But (the patient) already went through the complete eye check; the doctor know her condition… I don’t want to come and visit the doctor [again]… it will waste 2-3 hours.” (A21).
@@
When participants received an “I don’t know” response, they sought information from other sources, particularly the internet. For time-sensitive questions, participants reported using the internet to “double-check” the initial LLM-generated, unverified response due to delays in receiving an expert verified response. Four participants, accustomed to relying on the internet for their informational needs, reserved CataractBot exclusively for questions requiring expert opinions, such as “When can i do exercises as anulom vilom?” (P27). According to D4, without CataractBot, patients might have resorted to the internet and risked following inaccurate advice, given the unavailability of doctors for immediate consultations. CataractBot improves the likelihood of receiving accurate responses, as it uses a custom knowledge base.
6.2. Effect of Key Features on the Bot’s Usage
6.2.1. LLM-powered
Patients and attendants found the LLM-generated responses from the curated knowledge base were useful, and experts described them as “highly accurate”, noting that they minimized their workload for corrections. These instant responses saved time as information seekers did not have to constantly wait, either physically or over a call. Seven participants emphasized that traveling is time-consuming, and contacting busy hospitals over the phone is challenging due to high patient loads, as “the phone is always busy” (A4). Further, the robustness of the GPT-4 model enabled CataractBot to appropriately handle complex queries from participants with limited literacy, who struggled to formulate questions, consistent with prior research (Jain et al., 2018). For example, P18 asked “Tomorrow operation When should come No phone call received yet?”. CataractBot was able to address 45/65 such questions, of which 26 were asked using audio in an Indic language. Typing in Indic languages using the Latin script (e.g., instead of Devanagari script for Hindi) was rare, occurring in only three messages. In two instances, the messages mixed Hindi and English, and the bot responded accurately:
@@ 6.2.1. LLM-powered
For experts, the LLM enabled them to swiftly correct the bot’s responses using natural language, without worrying about grammatical or spelling errors. For example, in response to a query, D2 added the correction: “Btr to book an appt”. In other instances, experts also instructed the LLM to remove specific sentences, relying on it to generate appropriate responses from their feedback.
6.2.2. Experts-in-the-loop
Information seekers appreciated the novel feature of having human experts verify and update the auto-generated responses (Figure 1F), serving as the key reason for their trust in and engagement with the bot. For instance, “That’s why I gave my WhatsApp (number to onboard the bot), so I can talk directly to the doctor.” (A2). A12 added: “The trust only came when I saw the green tick mark… Before that, I was also in the question mark zone… thinking it’s a machine… But then somebody is checking it, confirming it… It’s actually trustworthy then.” Despite our concerns that the icons might be complex, participants found them self-explanatory:
Although the experts-in-the-loop feature increased trust in the final responses, it had a negative impact on perceived bot intelligence. Four participants noted that when an expert corrected CataractBot’s initial answer, their trust in the bot’s unverified answers reduced. For instance, “After some time, when the doctor said it is invalid, I was like, okay, should I even trust the bot?” (P19). In total, 12.8% (44) of the bot’s generated answers (excluding “I don’t know”) for medical and logistical queries were marked incorrect by the expert. This could be due to the limited custom knowledge base, the experts’ high standards for tone and structure of answers, or our prompting strategy favoring caution. Also, since participants knew that these messages were being reviewed by experts, four used the bot to send direct messages to them instead of asking questions (e.g., “Dr. Give me a call?” (P19)).
Experts appreciated the intermediary role of the bot, introducing a layer of separation from patients: “From a doctor’s standpoint, it is very good because it puts in a curtain between us.” (D4). Further analysis revealed that their suggested edits fell into five categories: adding missing information (76.8%), asking clarification questions (14.6%), making factual corrections (11.0%), asking the patient to visit the hospital (8.8%), and removing unnecessary information (3.7%). The length of expert corrections in characters varied significantly (t(91)=5.2, p¡0.05) with 152.4104.9 for medical questions and 50.942.6 for logistical questions.
Finally, experts did not have access to the final answer sent to information seekers after their suggested edits. This limited verification to a single round, and was designed to minimize their workload. However, this lack of transparency sometimes left experts uncertain if CataractBot had understood their corrections. Only the knowledge base update expert had access to the updated answer sent to the information seeker.
6.2.3. Multimodality
The 19 patients sent 89 messages (43 text, 26 audio, and 20 taps), while the 30 attendants sent 254 messages (163 text, 17 audio, and 74 taps) (Figure 3C). One-way ANOVAs on message types found statistical significance for attendants (F(2, 87)=13.74, p¡.001), but not for patients (F(2, 54)=0.77, p=0.46). Post-hoc pairwise comparisons, corrected with Holm’s sequential Bonferroni procedure, indicated statistically significant differences between ‘text’ and ‘audio’ (t(29)=4.8, p¡.01) and ‘text’ and ‘taps’ (t(29)=3.0, p¡0.01) for attendants. The minimal usage of audio by attendants could be due to demographic differences, with the majority being under 45 years and well-educated. Audio messages were predominantly used by older semi-literate patients. For instance, “I can’t read (or) write Hindi that well, so asking and listening to audio, I did that.” (P12).
Post-surgery limitations on screen use also indicated that speech input and audio output the preferred modality for patients, with 18 out of 26 audio questions by patients asked post-surgery. 31 participants used the related questions at least once. As P7 stated: “I used the audio message once. After that there was no need for me [to type or use audio]… I just kept selecting from the [related questions] options.” While they found the related questions feature convenient and frequently used it, there were instances of dissatisfaction when (in eight cases) the answers to those resulted in “I don’t know” responses. This issue arose because the CataractBot system uses an LLM call to suggest three questions based on the users’ last question, independent of the knowledge base. This approach aimed to uncover gaps in the knowledge base by not restricting questions to the curated knowledge base.
On the expert end, they could only interact with CataractBot using text. This design choice facilitated a “hygiene check” (D4) for thoroughness and precision in experts’ edits, aligning with previous findings (Li et al., 2023).
Patients mentioned utilizing the persistent information provided by CataractBot as a reference and sharing it with others undergoing cataract surgery. Also, as per D1, this information can help when discussing their surgery with experts outside of the Sankara Eye Hospital ecosystem. However, it also increased the sense of accountability among experts, making them “somewhat nervous”. They felt the need to be “very lawyer-like in our conversation because we don’t want it to come back and bite us tomorrow.” (D4).
6.2.4. Multilinguality
The majority of information seekers (29 out of 49) selected English, while 7 opted for Hindi, 6 Kannada, 5 Tamil, and 2 Telugu. None of the participants changed their initially selected language. Given the challenges of typing in Indic languages on smartphones (Joshi et al., 2011), our non-English participants predominantly relied on speech input and selecting from related questions. In total, patients who chose English typed 35 questions, asked 0 questions in speech, and tapped 6 related questions. In contrast, patients opting for Indic languages typed 8 questions, sent 26 audio questions and tapped 14 related questions in total. A similar trend was observed among attendants. Patients found value in local language interaction as input, helping them use the bot independently: “Without even asking me for help, he (patient) was able to use that bot as it supports Tamil.” (A13).
The audio queries in Indic languages were transcribed and translated to English using Azure AI services. However, this was not highly accurate due to the variability in Indic language dialect. This is a known issue and prior work (Jain et al., 2018) relied on human wizards to correct such errors. Experts faced challenges in “deciphering” such messages since they only received the English transcription. In the future, there should be a way to share the original audio query with the expert on demand.
6.2.5. Support for Multiple Stakeholders
Patients sent a total of 89 messages (4.75.4 message/patient), while attendants sent 254 messages (8.56.8 message/attendant), significantly more than patients, with t(47)=2.04, p¡0.05 (Figure 3C). During our interviews, we identified a few reasons for the patients to minimally use the bot, particularly related to demographics. Patients were typically older with lower levels of literacy and tech-literacy compared to their attendants These patients trusted the hospital to provide them with all the required information and did not actively seek information themselves. For instance:
“Dad didn’t try… I was the one thinking about it [the surgery]. He was like, “It’s OK! Not everyone needs to bother in the family… (to know) about what will be this, what will be that”. ” – A3.
Additionally, patients were advised to minimize screen use post-surgery, contributing to reduced bot interaction. Interestingly, patients sometimes offloaded the knowledge-seeking ‘work’ to their active attendants—e.g., “I have been using it. I told him [the patient] what answers I got. So maybe he did not feel the need (to use it).” (A20).
Moving on to experts, both the doctors and patient coordinators highlighted the significant role of escalation experts. Among the 309 verified medical and logistical answers, 62.5% were verified by the operating doctor/coordinator and the rest 37.5% by the escalation expert. Doctors mentioned having 2-3 days per week reserved for surgeries, and on these “operating days”, they rarely checked their phones during work hours, leading to more escalations. Despite their crucial role, escalation experts interacted differently from operating doctors and coordinators. Unlike the latter, who had face-to-face meetings with patients, escalation experts lacked contextual knowledge, limiting them to providing generic responses.
6.2.6. Knowledge Base